OpenStudy (anonymous):

Find the distance between the points (4,-3) and (-6,2).

OpenStudy (anonymous):

I got to : \[\sqrt{(-6-4)^2+(2+3)^2}\]

OpenStudy (anonymous):

And then I came out with \[\sqrt{65}\]

OpenStudy (cggurumanjunath):

http://prntscr.com/2kyomi

OpenStudy (cggurumanjunath):

sqrt (125)

OpenStudy (cggurumanjunath):

steps are correct but final answer is wrong @ilmc1

OpenStudy (anonymous):

Oh okay, how do you simplify \[\sqrt{125}\]

OpenStudy (cggurumanjunath):

125=5*5*5

OpenStudy (cggurumanjunath):

sqrt (125)=sqrt (5*5^2)=5*sqrt(5)

OpenStudy (anonymous):

Ohhh okay thanks :)

OpenStudy (cggurumanjunath):

:)